My wife and I exercise on a running course that is 2.25 miles long. If we both started at the same spot but traveled opppsite directions on the course and she walks 3 mph and I run 6mph, at what point in the course will we meet in distance and in time?
Answer by stanbon(75887)

My wife and I exercise on a running course that is 2.25 miles long. If we both started at the same spot but traveled opposite directions on the course and she walks 3 mph and I run 6mph, at what point in the course will we meet in distance and in time?
---------------------
Your DATA:
rate = 6 mph ; time = x hrs ; distance = rt = 6x miles
-----
Wife DATA:
rate = 3 mph ; time = x hrs ; distance = rt = 3x
----
Equation:
6x+3x = 2.25 miles
9x = 2.25 miles
x = 0.25 miles
---
Your distance: 6x = 6*0.25 = 1.5 miles
Wife's distance: 3x = 3*0.25 = 0.75 miles
--
Time = distance/rate = 1.5 miles/6 mph = 0.25 hrs
=====================================================
Cheers,
Stan H.
===========